# Dissipating quartets of excitations in a superconducting circuit

### v2

@@ Acknowledgments
$V_{\mathrm{chain}}(\psi)=-NE^{0}_{J}\cos(\psi/N)\simeq-\frac{E^{0}_{J}}{2N}\psi^{2}+\frac{E^{0}_{J}}{24N^{3}}\psi^{4}$ (S14)
@@
S2.4 Miscalibration of the ATS flux biases
While the ATS should in principle be driven with symmetric fluxes to activate the desired four-to-one photon exchange (see Sec. I), we actually drive the circuit with an unknown combination of common and differential flux drives (see Sec. S1.2). Dynamical renormalization of the modes non-linearity may then arise.
To understand this phenomenon, we consider the ideal circuit Hamiltonian (S3) (other sources of non-linearity neglected for clarity), at the ATS operating point and in presence of flux drives biasing the loops both symmetrically (reduced amplitude $\epsilon_{c}(t)$) and antisymmetrically (reduced amplitude $-\epsilon_{d}(t)$). In the rotating frame, it reads
${\bf H}_{cd}(t)=E_{\Sigma}\sin\big{(}\epsilon_{c}(t)\big{)}\sin\big{(}\bm{\varphi}(t)+\epsilon_{d}(t)\big{)}$ (S15)
where $\bm{\varphi}(t)=\sum_{j=a,b,r}\varphi^{ZPF}_{j}({\bf a}_{j}e^{-i\omega_{j}t}+{\bf a}_{i}^{\dagger}e^{+i\omega_{j}t})$. Note that we expressed the Hamiltonian in a gauge where the differential drive appears inside the sine term, which may lead to the apparition of an effective charge drive in the Hamiltonian [85]. The amplitude of this term depends on the exact magnetic field distribution with respect to the capacitive structures in the circuit and is hard to predict a priori. Since it does not yield resonant terms in the RWA, we neglect it in our analysis for simplicity. However, we point out that this drive—along with an effective flux drive stemming from the first-order term in the sine expansion—displaces the states of the modes, yielding new dynamical corrections to the modes non-linearity at the second order of the RWA.
@@
${\bf H}_{cd}(t)=E_{\Sigma}~{}\epsilon_{c}(t)\sin\big{(}\bm{\varphi}(t)\big{)}+E_{\Sigma}~{}\epsilon_{c}(t)\epsilon_{d}(t)\cos\big{(}\bm{\varphi}(t)\big{)}$ (S16)
@@
${\bf H}_{cd,RWA}=\overline{{\bf H}_{cd}(t)}=2\Re\big{(}\xi_{c}\xi^{\ast}_{d}\big{)}E_{\Sigma}~{}\overline{\cos\big{(}\bm{\varphi}(t)\big{)}}$ (S17)
This effective Hamiltonian is analogous to the non-linear term in the Hamiltonian (S9) deriving from imbalance of the SQUID. In practice, it slightly renormalizes the modes non-linearity, which we do not account for in our simulations (as detailed in Sec. S4.2, the system dynamics under pump does not depend on the exact value of the non-linearity).
We also note that this term is resonant irrespective of the pump frequency. Thus, it should be possible to apply an off-resonant pump ($\omega_{p}\neq 4\omega_{a}-\omega_{b}$) to dynamically cancel the non-linearity stemming from imbalance of the SQUID. One simply adjusts the symmetric and antisymmetric drive amplitudes and phases such that $2\Re\big{(}\xi_{c}\xi^{\ast}_{d}\big{)}=E_{\Delta}$, where $E_{\Delta}$ is the difference of Josephson energy for the two junctions. However, given that the above analysis is only valid in the limit $|\epsilon_{c}|,|\epsilon_{d}|\ll 1$, this method may only prove useful in the limit $E_{\Delta}\ll E_{\Sigma}$. Finally, we note that the effective Hamiltonian (S17) also induces a shift of the modes frequency. This mechanism is expected to contribute to the controlled frequency shift (induced by an off-resonant pump) that we apply to avoid hybridization of our system with two-level systems (see Sec. S5).
S3 Details of the calibration and measurement sequences
S3.1 Photon-number resolving measurement
Figure S6: Reference signals for the photon-number resolving measurement. In order to maximize SNR, we acquire reference signals $r_{n}^{ON}(t)$ and $r_{n}^{OFF}(t)$ after preparing the state $|n\rangle_{a}$ and applying (ON) or not (OFF) a pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition. We use each signal as an averaging envelope for subsequent measurement of the occupation of $|n\rangle_{a}$ for an unknown state (Eq. (S18)). The non-zero signal at long time for $n=0$ is attributed to residual thermal occupation of state $|n=1\rangle_{1}$ at equilibrium. The maximum of each signal is expected to be delayed when $n$ increases, which is clearly visible for $n=4$ (the initial drop is not understood). The pulses duration and amplitude (referenced by the voltage applied to the mixer generating the pulse) are 0.3 $\mu$s / 0.4 V, 1 $\mu$s / 0.15 V, 1 $\mu$s / 0.15 V and 0.5 $\mu$s / 0.1 V for $n=1,~{}2,~{}3$ and $4$ respectively. These characteristics are only optimized to maximize signal for $n=4$, which explains a higher signal than for $n=2,3$. Preparation of $|n=2\rangle_{a}$, which relies on two successive $\pi$-pulses on single-photon transitions is less efficient than the preparation of $|n=3\rangle_{a}$ (single three-photon $\pi$-pulse), explaining a weaker signal.
@@
While the memory mode vacuum state $|0\rangle_{a}$ exerts a dispersive pull onto the readout mode significantly different from the pull exerted by excited states, the differential pull exerted by one excited state and the next tails off rapidly. Indeed, the dispersive pulls stem from the $\cos(2\bm{\varphi})$ contribution in the circuit static Hamiltonian (see Sec. S2) and follow the same trend as the memory mode anharmonicity (see Fig. 2). This makes excited states hard to distinguish directly. We circumvent this shortcoming by applying (ON) or not (OFF) a $\pi$-pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition before probing the readout mode on resonance with a long (40 $\mu$s) pulse. The reflected field is demodulated and its quadratures averaged to obtain the complex signals $s_{n}^{ON}(t)$ and $s_{n}^{OFF}(t)$. Since the dynamics of the decay from $|n+3\rangle_{a}$ to the vacuum state is delayed compared to the decay from $|n\rangle_{a}$, $s_{n}^{ON}(t)$ detected after the application of the $\pi$-pulse is distorted and time-shifted with respect to the signal $s_{n}^{OFF}(t)$ detected in absence of the pulse.
In order to maximize the measurement signal-to-noise ratio (SNR), we first prepare $|n\rangle_{a}$ with a sequence of $\pi$-pulses on single or three-photon transitions, and record an averaged reference signal $r_{n}^{ON}(t)$ and $r_{n}^{OFF}(t)$ (see Fig. S6). Subsequently, when measuring an unknown state of the memory, we integrate the detected differential signal with an envelope provided by these reference signals:
$S_{n}=\frac{1}{T^{1/2}_{\mathrm{bin}}}\dfrac{\int_{t=0}^{T}\Re\Big{(}\big{(}s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big{)}\big{(}r_{n}^{ON}(t)-r_{n}^{OFF}(t)\big{)}^{\ast}\Big{)}\mathrm{d}t}{\Big{(}\int_{t=0}^{T}|r_{n}^{ON}(t)-r_{n}^{OFF}(t)|^{2}~{}\mathrm{d}t\Big{)}^{1/2}}$ (S18)
@@
In order to maximize SNR while ensuring that the $\pi$-pulse does not impact the system if only states $|k\rangle_{a}$ with $k<n$ are occupied, we need to adjust finely its amplitude and duration. Intuitively, we want to minimize the pulse duration while ensuring that it is sufficiently weak and well-resolved in frequency not to excite the $|n-1\rangle_{a}\leftrightarrow|n+2\rangle_{a}$ transition. To this end, we prepare in separate experiments the state $|n\rangle_{a}$ and the state $|n-1\rangle_{a}$ and vary the pulse amplitude and duration (square waveform with smoothed edges). We then select the characteristics yielding the highest signal when $|n\rangle_{a}$ is prepared while yielding a negligible signal when $|n-1\rangle_{a}$ is prepared (see Fig. S7). Note that after this calibration, we could again acquire reference signals for the integration of $S_{n}$ using the optimal $\pi$-pulse characteristics. This was done for $S_{4}$, but not for $S_{n}$ with $n\leq 3$ (we have verified that the reference signal shape does not depend significantly on the $\pi$-pulse characteristics as long as it does not excite unwanted transitions).
@@
Figure S7: Optimization of the $\pi$-pulses parameters for the photon-number resolving measurement. The measurement relies on the differential signal obtained when exciting or not the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition. Holding the exciting pulse duration constant, we vary its amplitude and record the differential signal $S_{n}$ after preparing $|n\rangle_{a}$ (plain lines) or $|n-1\rangle_{a}$ (dashed lines). We select the amplitude yielding the strongest signal in the former case while yielding a negligible signal in the latter case (vertical lines). We repeat this calibration for various pulse durations and select the optimal amplitude combination of amplitude and duration.
S3.2 Tuning the pump frequency
@@
Figure S8: Pump frequency calibration. The pump and buffer frequencies dressed by the pump itself are calibrated iteratively (see text). The buffer frequency $\omega_{b_{1}}$ found at step 1 and $\omega_{b_{3}}$ found at step 3 are slightly mismatched at low pump amplitude. This is attributed to inaccuracy of the estimate made at step 3, in which the reflection was not properly fitted but only smoothed aggressively before locating the minimum of the signal magnitude. This mismatch then propagates to the calibrated pump frequency (step 2 vs step 4) as the pump resonance condition reads $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n-1}}$ where $\omega_{b_{n-1}}$ is the frequency at which the buffer is driven (found at the previous step) and $\omega_{a_{n}}$ is the memory frequency dressed by the pump at $\omega_{p_{n}}$.
@@
In Fig. 3, we report the decay of population in state $|4\rangle_{a}$ and the buffer reflection spectrum as the amplitude of the pump addressing the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition is varied (encoded in color). For each pump amplitude, the pump frequency needs to be adjusted as the modes resonance frequency gets shifted through by the pump itself (shifting mechanisms include—but are not limited to—miscalibration of the pump symmetry described in Sec. S2 and direct Stark shift owing to the modes non-linearity). This shift depends non trivially on the pump frequency, mostly due to the uncalibrated dispersion of the cryostat input lines. For a given pump amplitude, the pump frequency is adjusted following four steps.
@@
    Again as in step (1), we hold the pump at $\omega_{p_{2}}$ and probe the buffer to find its resonance frequency $\omega_{b_{3}}$. Note that the buffer spectra acquired at this step are the ones reported in Fig. 3 and in Fig. S10. For the largest pump amplitudes, the buffer line splits owing to the hybridization of $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$. Fitting to the full two-mode model presented in Sec. S4.2 was not attempted at this step, so that the buffer resonance frequency was simply extracted by smoothing each spectrum over a wide moving window and locating the minimum of the smoothed signal magnitude.
@@
One should iterate further the procedure until the resonance condition $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n}}$ is met self-consistently. In practice, the variation of the pump frequency $|\omega_{p_{2}}-\omega_{p_{4}}|$ between step (2) and (4) was deemed small enough (less than 0.6 MHz at all considered pump amplitudes) that the buffer frequency $\omega_{b_{4}}$ dressed by the the pump at $\omega_{p_{4}}$ would be near identical the frequency $\omega_{b_{3}}$ measured under the pump at $\omega_{p_{2}}$. This assumption, together with the rather coarse estimate of $\omega_{b_{1}}$ in step (3) explains a small detuning between the predictions from our model (plain lines in Fig. S10) to the recorded spectra (dots).
@@
S3.3 Decay rate of low energy states under pump
@@
As a sanity check of our experiment, we record, for the same range of pump amplitudes activating the four-photon dissipation, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ (see Fig. S9). For all states, we observe a slight increase of the decay rate induced by the pump, which is not captured by our model. For the states $|2\rangle_{a}$ and $|3\rangle_{a}$, we also record an initial sharp drop of the corresponding signals $S_{2}$ and $S_{3}$. This drop is not understood, but does not seem directly induced by the pump as it is also visible for a null pump amplitude (dark blue). We thus rather suspect an effect caused by the proximity of the memory preparation pulses and of the final pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition (used in the measurement) for the shortest wait times. All signals are fitted with exponential laws and the fitted decay rates are reported in Fig. 4.
@@
Figure S9: Decay rate of the low-energy states under pump. The time decay under pump of population in states $|1\rangle_{a}$ (top), $|2\rangle_{a}$ (middle) and $|3\rangle_{a}$ (bottom) is recorded (filled circles) and fitted with an exponential law (lines). For each state $|n\rangle_{a}$, we fit a single scaling factor for the differential signal $S_{n}$, close to the value calibrated in Fig. S7 (the slightly lower values recorded here, in particular for $S_{3}$, are attributed to drifts during the $\sim 12$ hour acquisition time). The stationary value of the exponential law is fixed to zero, except for $S_{1}$ where we leave it as a free fit parameter to account for thermal excitations of the memory at equilibrium (see Fig. S6). The initial drop of $S_{2}$ and $S_{3}$ is not understood but does not seem directly related to the pump as it is visible at all pump amplitudes.
@@
Figure S10: Buffer reflection spectrum under pump. We represent the amplitude of the field reflected off the buffer as in Fig. 3c (top, curves offset for readability), together with the phase of the reflected signal (middle) and a representation in the Fresnel plane (bottom). The spectra are well reproduced by simulations based on the effective Hamiltonian (S35) (lines).
S4 Numerical simulations and fitting procedures
S4.1 Frequency maps

### v3

@@ B.3 Non-linearity induced by the superinductors
$V_{\mathrm{chain}}(\psi)=-NE^{0}_{J}\cos(\psi/N)\simeq-\frac{E^{0}_{J}}{2N}\psi^{2}+\frac{E^{0}_{J}}{24N^{3}}\psi^{4}$ (25)
@@
B.4 Miscalibration of the ATS flux biases
While the ATS should in principle be driven with symmetric fluxes to activate the desired four-to-one photon exchange (see Sec. I), we actually drive the circuit with an unknown combination of common and differential flux drives (see Sec. A.2). Dynamical renormalization of the modes non-linearity may then arise.
To understand this phenomenon, we consider the ideal circuit Hamiltonian (8) (other sources of non-linearity neglected for clarity), at the ATS operating point and in presence of flux drives biasing the loops both symmetrically (reduced amplitude $\epsilon_{c}(t)$) and antisymmetrically (reduced amplitude $\epsilon_{d}(t)$). In the rotating frame, it reads
${\bf H}_{cd}(t)=E_{\Sigma}\sin\big(\epsilon_{c}(t)\big)\sin\big(\boldsymbol{\varphi}(t)+\epsilon_{d}(t)\big)$ (26)
where $\boldsymbol{\varphi}(t)=\sum_{j=a,b,r}\varphi^{ZPF}_{j}({\bf a}_{j}e^{-i\omega_{j}t}+{\bf a}_{i}^{\dagger}e^{+i\omega_{j}t})$. Note that we expressed the Hamiltonian in a gauge where the differential drive appears inside the sine term, which may lead to the apparition of an effective charge drive in the Hamiltonian [86]. The amplitude of this term depends on the exact magnetic field distribution with respect to the capacitive structures in the circuit and is hard to predict a priori. Since it does not yield resonant terms in the RWA, we neglect it in our analysis for simplicity. However, we point out that this drive—along with an effective flux drive stemming from the first-order term in the sine expansion—displaces the states of the modes, yielding new dynamical corrections to the modes non-linearity at the second order of the RWA.
@@ B.4 Miscalibration of the ATS flux biases
${\bf H}_{cd}(t)=E_{\Sigma}~\epsilon_{c}(t)\sin\big(\boldsymbol{\varphi}(t)\big)+E_{\Sigma}~\epsilon_{c}(t)\epsilon_{d}(t)\cos\big(\boldsymbol{\varphi}(t)\big)$ (27)
@@
${\bf H}_{cd,RWA}=\overline{{\bf H}_{cd}(t)}=2\Re\big(\xi_{c}\xi^{\ast}_{d}\big)E_{\Sigma}~\overline{\cos\big(\boldsymbol{\varphi}(t)\big)}$ (28)
This effective Hamiltonian is analogous to the non-linear term in the Hamiltonian (14) deriving from imbalance of the SQUID. In practice, it slightly renormalizes the modes non-linearity, which we do not account for in our simulations (as detailed in Sec. D.2, the system dynamics under pump does not depend on the exact value of the non-linearity).
We also note that this term is resonant irrespective of the pump frequency. Thus, it should be possible to apply an off-resonant pump ($\omega_{p}\neq 4\omega_{a}-\omega_{b}$) to dynamically cancel the non-linearity stemming from imbalance of the SQUID. One simply adjusts the symmetric and antisymmetric drive amplitudes and phases such that $2\Re\big(\xi_{c}\xi^{\ast}_{d}\big)=E_{\Delta}$, where $E_{\Delta}$ is the difference of Josephson energy for the two junctions. However, given that the above analysis is only valid in the limit $|\epsilon_{c}|,|\epsilon_{d}|\ll 1$, this method may only prove useful in the limit $E_{\Delta}\ll E_{\Sigma}$. Finally, we note that the effective Hamiltonian (28) also induces a shift of the modes frequency. This mechanism is expected to contribute to the controlled frequency shift (induced by an off-resonant pump) that we apply to avoid hybridization of our system with two-level systems (see Sec. E).
Appendix C Details of the calibration and measurement sequences
C.1 Photon-number resolving measurement
[…]
@@ C.1 Photon-number resolving measurement
[…]
In order to maximize the measurement signal-to-noise ratio (SNR), we first prepare $|n\rangle_{a}$ with a sequence of $\pi$-pulses on single or three-photon transitions, and record an averaged reference signal $r_{n}^{ON}(t)$ and $r_{n}^{OFF}(t)$ (see Fig. 10). Subsequently, when measuring an unknown state of the memory, we integrate the detected differential signal with an envelope provided by these reference signals:
$S_{n}=\frac{1}{T^{1/2}_{\mathrm{bin}}}\dfrac{\int_{t=0}^{T}\Re\Big(\big(s_{n}^{ON}(t)-s_{n}^{OFF}(t)\big)\big(r_{n}^{ON}(t)-r_{n}^{OFF}(t)\big)^{\ast}\Big)\mathrm{d}t}{\Big(\int_{t=0}^{T}|r_{n}^{ON}(t)-r_{n}^{OFF}(t)|^{2}~\mathrm{d}t\Big)^{1/2}}$ (29)
@@
In order to maximize SNR while ensuring that the $\pi$-pulse does not impact the system if only states $|k\rangle_{a}$ with $k<n$ are occupied, we need to adjust finely its amplitude and duration. Intuitively, we want to minimize the pulse duration while ensuring that it is sufficiently weak and well-resolved in frequency not to excite the $|n-1\rangle_{a}\leftrightarrow|n+2\rangle_{a}$ transition. To this end, we prepare in separate experiments the state $|n\rangle_{a}$ and the state $|n-1\rangle_{a}$ and vary the pulse amplitude and duration (square waveform with smoothed edges). We then select the characteristics yielding the highest signal when $|n\rangle_{a}$ is prepared while yielding a negligible signal when $|n-1\rangle_{a}$ is prepared (see Fig. 11). Note that after this calibration, we could again acquire reference signals for the integration of $S_{n}$ using the optimal $\pi$-pulse characteristics. This was done for $S_{4}$, but not for $S_{n}$ with $n\leq 3$ (we have verified that the reference signal shape does not depend significantly on the $\pi$-pulse characteristics as long as it does not excite unwanted transitions).
@@
Figure 11: Optimization of the $\pi$-pulses parameters for the photon-number resolving measurement. The measurement relies on the differential signal obtained when exciting or not the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition. Holding the exciting pulse duration constant, we vary its amplitude and record the differential signal $S_{n}$ after preparing $|n\rangle_{a}$ (plain lines) or $|n-1\rangle_{a}$ (dashed lines). We select the amplitude yielding the strongest signal in the former case while yielding a negligible signal in the latter case (vertical lines). We repeat this calibration for various pulse durations and select the optimal amplitude combination of amplitude and duration.
C.2 Tuning the pump frequency
@@
Figure 12: Pump frequency calibration. The pump and buffer frequencies dressed by the pump itself are calibrated iteratively (see text). The buffer frequency $\omega_{b_{1}}$ found at step 1 and $\omega_{b_{3}}$ found at step 3 are slightly mismatched at low pump amplitude. This is attributed to inaccuracy of the estimate made at step 3, in which the reflection was not properly fitted but only smoothed aggressively before locating the minimum of the signal magnitude. This mismatch then propagates to the calibrated pump frequency (step 2 vs step 4) as the pump resonance condition reads $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n-1}}$ where $\omega_{b_{n-1}}$ is the frequency at which the buffer is driven (found at the previous step) and $\omega_{a_{n}}$ is the memory frequency dressed by the pump at $\omega_{p_{n}}$.
@@
In Fig. 3, we report the decay of population in state $|4\rangle_{a}$ and the buffer reflection spectrum as the amplitude of the pump addressing the $|4\rangle_{a}|0\rangle_{b}\leftrightarrow|0\rangle_{a}|1\rangle_{b}$ transition is varied (encoded in color). For each pump amplitude, the pump frequency needs to be adjusted as the modes resonance frequency gets shifted through by the pump itself (shifting mechanisms include—but are not limited to—miscalibration of the pump symmetry described in Sec. B and direct Stark shift owing to the modes non-linearity). This shift depends non trivially on the pump frequency, mostly due to the uncalibrated dispersion of the cryostat input lines. For a given pump amplitude, the pump frequency is adjusted following four steps.
@@
    Again as in step (1), we hold the pump at $\omega_{p_{2}}$ and probe the buffer to find its resonance frequency $\omega_{b_{3}}$. Note that the buffer spectra acquired at this step are the ones reported in Fig. 3 and in Fig. 14. For the largest pump amplitudes, the buffer line splits owing to the hybridization of $|4\rangle_{a}|0\rangle_{b}$ and $|0\rangle_{a}|1\rangle_{b}$. Fitting to the full two-mode model presented in Sec. D.2 was not attempted at this step, so that the buffer resonance frequency was simply extracted by smoothing each spectrum over a wide moving window and locating the minimum of the smoothed signal magnitude.
@@
One should iterate further the procedure until the resonance condition $\omega_{p_{n}}=4\omega_{a_{n}}-\omega_{b_{n}}$ is met self-consistently. In practice, the variation of the pump frequency $|\omega_{p_{2}}-\omega_{p_{4}}|$ between step (2) and (4) was deemed small enough (less than 0.6 MHz at all considered pump amplitudes) that the buffer frequency $\omega_{b_{4}}$ dressed by the the pump at $\omega_{p_{4}}$ would be near identical the frequency $\omega_{b_{3}}$ measured under the pump at $\omega_{p_{2}}$. This assumption, together with the rather coarse estimate of $\omega_{b_{1}}$ in step (3) explains a small detuning between the predictions from our model (plain lines in Fig. 14) to the recorded spectra (dots).
@@
C.3 Decay rate of low energy states under pump
@@
As a sanity check of our experiment, we record, for the same range of pump amplitudes activating the four-photon dissipation, the time-decay of populations prepared in states $|n\rangle_{a}$ for $n<4$ (see Fig. 13). For all states, we observe a slight increase of the decay rate induced by the pump, which is not captured by our model. For the states $|2\rangle_{a}$ and $|3\rangle_{a}$, we also record an initial sharp drop of the corresponding signals $S_{2}$ and $S_{3}$. This drop is not understood, but does not seem directly induced by the pump as it is also visible for a null pump amplitude (dark blue). We thus rather suspect an effect caused by the proximity of the memory preparation pulses and of the final pulse on the $|n\rangle_{a}\leftrightarrow|n+3\rangle_{a}$ transition (used in the measurement) for the shortest wait times. All signals are fitted with exponential laws and the fitted decay rates are reported in Fig. 4.
@@
Figure 13: Decay rate of the low-energy states under pump. The time decay under pump of population in states $|1\rangle_{a}$ (top), $|2\rangle_{a}$ (middle) and $|3\rangle_{a}$ (bottom) is recorded (filled circles) and fitted with an exponential law (lines). For each state $|n\rangle_{a}$, we fit a single scaling factor for the differential signal $S_{n}$, close to the value calibrated in Fig. 11 (the slightly lower values recorded here, in particular for $S_{3}$, are attributed to drifts during the $\sim 12$ hour acquisition time). The stationary value of the exponential law is fixed to zero, except for $S_{1}$ where we leave it as a free fit parameter to account for thermal excitations of the memory at equilibrium (see Fig. 10). The initial drop of $S_{2}$ and $S_{3}$ is not understood but does not seem directly related to the pump as it is visible at all pump amplitudes.
@@
Figure 14: Buffer reflection spectrum under pump. We represent the amplitude of the field reflected off the buffer as in Fig. 3c (top, curves offset for readability), together with the phase of the reflected signal (middle) and a representation in the Fresnel plane (bottom). The spectra are well reproduced by simulations based on the effective Hamiltonian (46) (lines).
Appendix D Numerical simulations and fitting procedures
D.1 Frequency maps